# Automating Continual Learning

### v1

@@ Abstract
General-purpose learning systems should improve themselves in open-ended fashion in ever-changing environments. Conventional learning algorithms for neural networks, however, suffer from catastrophic forgetting (CF)—previously acquired skills are forgotten when a new task is learned. Instead of hand-crafting new algorithms for avoiding CF, we propose Automated Continual Learning (ACL) to train self-referential neural networks to meta-learn their own in-context continual (meta-)learning algorithms. ACL encodes all desiderata—good performance on both old and new tasks—into its meta-learning objectives. Our experiments demonstrate that ACL effectively solves “in-context catastrophic forgetting”; our ACL-learned algorithms outperform hand-crafted ones, e.g., on the Split-MNIST benchmark in the replay-free setting, and enables continual learning of diverse tasks consisting of multiple few-shot and standard image classification datasets. [^1] Our code is public: https://github.com/IDSIA/automated-cl. ${}^{\dagger}$ork done at IDSIA.
@@ 1 Introduction
Here, instead of hand-crafting learning algorithms for continual learning, we train self-referential neural networks (Schmidhuber, 1992a; 1987) to meta-learn their own “in-context” continual learning algorithms. We train them through gradient descent on learning objectives that reflect desiderata for continual learning algorithms—good performance on both old and new tasks, including forward and backward transfer. In fact, by extending the standard settings of few-shot or meta-learning based on sequence-processing NNs (Hochreiter et al., 2001; Younger et al., 1999; Cotter & Conwell, 1991; 1990; Mishra et al., 2018), the continual learning problem can also be formulated as a long-span sequence processing task (Irie et al., 2022c). Corresponding sequences can be obtained by concatenating multiple few-shot/meta-learning sub-sequences, where each sub-sequence consists of input/target examples corresponding to the task to be learned in-context. As we’ll see in Sec. 3, this setting also allows us to seamlessly express classic desiderata for continual learning (knowledge preservation, forward/backward transfer, etc.) as part of objective functions of the meta-learner.
Once formulated as such a sequence-learning task, we let gradient descent search for CL algorithms achieving the desired CL behaviors in the program space of NN weights. In principle, all typical challenges of CL—such as the stability-plasticity dilemma (Grossberg, 1982)—are automatically discovered and handled by the gradient-based program search process. Once trained, CL is automated through recursive self-modification dynamics of the trained NN, without requiring any human intervention such as adding extra regularization or setting hyper-parameters for continual learning. Therefore, we call our method, Automated Continual Learning (ACL).
@@ 2 Background
Here we briefly review some background concepts that are essential for describing our method in Sec. 3: continual learning and its desiderata (Sec. 2.1), few-shot/meta-learning via sequence processing (Sec. 2.2), and linear Transformer/fast weight programmer architectures (Sec. 2.3) that are foundations of the self-referential neural network we use in our experiments.
@@ 2.1 Continual Learning
Continual or lifelong learning is a special form of multi-task learning (Thrun, 1998; Caruana, 1997). In conventional multi-task learning scenarios, all datasets for different tasks are available at once, and NNs are trained jointly on all of them (e.g., training batches mix examples from different datasets without any particular order). In contrast, in continual learning settings, presentation of tasks/datasets is sequential; an NN training process only has access to one task/dataset at a time. When a dataset for a new task becomes available, we lose access to the training examples from the previous task, and so on.
@@
Classic desiderata for a CL system (see, e.g., Lopez-Paz & Ranzato (2017); Veniat et al. (2021)) are typically summarized as good performance on three metrics: classification accuracies on each dataset (or their average), backward transfer (i.e., impact of learning a new task on the model’s performance on previous tasks; e.g., catastrophic forgetting is a negative backward transfer), and forward transfer (impact of learning a task for the model’s performance on a future task). From a broader perspective of meta-learning systems, we may also measure other effects such as learning acceleration (i.e., whether the system leverages previous learning experiences to accelerate future learning); here we only briefly discuss this as our primary focus remains the classic CL metrics above.
@@ 2.2 Few-shot/meta-learning via Sequence Learning
In Sec. 3, we’ll formulate continual learning as a long-span sequence processing task. This is a direct extension of the classic few-shot/meta learning formulated as a sequence learning problem. In fact, since the seminal works by Hochreiter et al. (2001); Younger et al. (1999); Cotter & Conwell (1991; 1990) (see also Naik & Mammone (1992)), many sequence processing neural networks (see, e.g., Bosc (2015); Santoro et al. (2016); Duan et al. (2016); Wang et al. (2017); Munkhdalai & Yu (2017); Munkhdalai & Trischler (2018); Miconi et al. (2018; 2019); Munkhdalai et al. (2019); Kirsch & Schmidhuber (2021); Sandler et al. (2021); Huisman et al. (2023) including “quadratic” Transformers (Vaswani et al., 2017) as in Mishra et al. (2018)) have been trained as a meta-learner (Schmidhuber, 1987; 1992a) that learn by observing sequences of training examples (i.e., pairs of inputs and their labels) in-context. Here we briefly review such a formulation.
@@
Let $d$, $N$, $K$, $P$ be positive integers. In sequential $N$-way $K$-shot classification settings, a sequence processing NN with a parameter vector $\theta\in\mathbb{R}^{P}$ observes a pair (${\bm{x}}_{t}$, $y_{t}$) where ${\bm{x}}_{t}\in\mathbb{R}^{d}$ is the input and $y_{t}\in\{1,...,N\}$ is its label at each step $t\in\{1,...,N\cdot K\}$, corresponding to $K$ examples for each one of $N$ classes. After the presentation of these $N\cdot K$ examples (often called the support set), one extra input ${\bm{x}}\in\mathbb{R}^{d}$ (often called the query) is fed to the model without its true label but an “unknown label” token $\varnothing$ (number of input labels accepted by the model is thus $N+1$). The model is trained to predict its true label, i.e., the parameters of the model $\theta$ are optimized to maximize the probability $p(y|({\bm{x}}_{1},y_{1}),...,({\bm{x}}_{N\cdot K},y_{N\cdot K}),({\bm{x}},\varnothing);\theta)$ of the correct label $y\in\{1,...,N\}$ of the input query ${\bm{x}}$. Since class-to-label associations are randomized and unique to each sequence ($({\bm{x}}_{1},y_{1}),...,({\bm{x}}_{N\cdot K},y_{N\cdot K}),({\bm{x}},\varnothing)$), each such a sequence represents a new (few-shot or meta) learning experience to train the model. To be more specific, this is the synchronous label setting of Mishra et al. (2018) where the learning phase (observing examples $({\bm{x}}_{1},y_{1}),...,({\bm{x}}_{N\cdot K},y_{N\cdot K})$) is separated from the prediction phase (predicting label $y$ given $({\bm{x}},\varnothing)$). We opt for this variant in our experiments as we empirically find this (at least in our specific settings) more stable than the delayed label setting (Hochreiter et al., 2001) where the model has to make a prediction for every input, and the label is fed to the model with a delay of one time step.
2.3 Self-Referential Weight Matrices or Recursive Linear Transformers
@@
Our method (Sec. 3) can be applied to any sequence-processing NN architectures in principle. Nevertheless, certain architectures naturally fit better to parameterize a self-improving continual learner. Here we use the modern self-referential weight matrix (SRWM; Irie et al. (2022c; 2023)) to build a generic self-modifying NN. An SRWM is a weight matrix that sequentially modifies itself as a response to a stream of input observations (Schmidhuber, 1992a; 1993). The modern SRWM belongs to the family of linear Transformers (LTs) a.k.a. Fast Weight Programmers (FWPs; Schmidhuber (1991; 1992b); Katharopoulos et al. (2020); Choromanski et al. (2021); Peng et al. (2021); Schlag et al. (2021); Irie et al. (2021a)). Linear Transformers and FWPs are an important class of the now popular Transformers (Vaswani et al., 2017): unlike the standard ones whose computational requirements grow quadratically and whose state size grows linearly with the context length, LTs/FWPs’ complexity is linear and the state size is constant w.r.t. sequence length (like in the standard RNNs). This is an important property for in-context CL, since, conceptually, we want such a CL system to continue to learn for an arbitrarily long, lifelong time span. Moreover, the duality between linear attention and FWPs (Schlag et al., 2021)—and likewise, between linear attention and gradient descent-trained linear layers (Irie et al., 2022a; Aizerman et al., 1964)—have played a key role in certain theoretical analyses of in-context learning capabilities of Transformers (von Oswald et al., 2023a; Dai et al., 2023).
@@
The dynamics of an SRWM (Irie et al., 2022c) are described as follows. Let $d_{\text{in}}$, $d_{\text{out}}$, $t$ be positive integers, and $\otimes$ denote outer product. At each time step $t$, an SRWM ${\bm{W}}_{t-1}\in\mathbb{R}^{(d_{\text{out}}+2*d_{\text{in}}+1)\times d_{\text{in}}}$ observes an input ${\bm{x}}_{t}\in\mathbb{R}^{d_{\text{in}}}$, and outputs ${\bm{y}}_{t}\in\mathbb{R}^{d_{\text{out}}}$, while also updating itself to ${\bm{W}}_{t}$:
@@
where ${\bm{v}}_{t},\bar{{\bm{v}}}_{t}\in\mathbb{R}^{(d_{\text{out}}+2*d_{\text{in}}+1)}$ are value vectors, ${\bm{q}}_{t}\in\mathbb{R}^{d_{\text{in}}}$ and ${\bm{k}}_{t}\in\mathbb{R}^{d_{\text{in}}}$ are query and key vectors, and $\sigma(\beta_{t})\in\mathbb{R}$ is the learning rate. $\sigma$ and $\phi$ denote sigmoid and softmax functions respectively. $\phi$ is typically also applied to ${\bm{x}}_{t}$ in Eq. 1; here we follow Irie et al. (2022c)’s few-shot image classification setting, and use the variant without it. Eq. 3 corresponds to a rank-one update of the SRWM, from ${\bm{W}}_{t-1}$ to ${\bm{W}}_{t}$, through the delta learning rule (Widrow & Hoff, 1960; Schlag et al., 2021) where the self-generated patterns, ${\bm{v}}_{t}$, $\phi({\bm{k}}_{t})$, and $\sigma(\beta_{t})$, play the role of target, input, and learning rate of the learning rule respectively. The delta rule in FWPs is typically reported to be crucial broadly across many practical tasks (Schlag et al., 2021; Irie et al., 2021a; 2022b; Irie & Schmidhuber, 2023b).
The initial weight matrix ${\bm{W}}_{0}$ is the only trainable parameters of this layer, that encodes the initial self-modification algorithm. In practice, we use the layer above as a direct replacement to the self-attention layer in the Transformer architecture (Vaswani et al., 2017); we also use the multi-head version of the SRWM computation above. For further details, we refer to Irie et al. (2022c).
@@ 3 Method
Task Formulation. We formulate continual learning as a long-span sequence learning task. Let $D$, $N$, $K$, $L$ denote positive integers. Consider two $N$-way classification tasks ${\mathbf{A}}$ and ${\mathbf{B}}$ to be learned sequentially (as we’ll see, this can also be straightforwardly extended to three or more tasks; also the formulation here applies to both “meta-training” and “meta-test” phases, see Appendix A.1 for more on the terminology). We denote the respective training datasets as $\mathcal{A}$ and $\mathcal{B}$, and test sets as $\mathcal{A^{\prime}}$ and $\mathcal{B^{\prime}}$. We assume that each datapoint in these datasets consists of one input feature ${\bm{x}}\in\mathbb{R}^{D}$ of dimension $D$ (generically denoted as vector ${\bm{x}}$, but it is an image in all our experiments) and one label $y\in\{1,...,N\}$ denoting one out of $N$ classes. We consider two sequences of $L$ training examples $\left(({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L})\right)$ and $\left(({\bm{x}}^{\mathcal{B}}_{1},y^{\mathcal{B}}_{1}),...,({\bm{x}}^{\mathcal{B}}_{L},y^{\mathcal{B}}_{L})\right)$ sampled from the respective training sets $\mathcal{A}$ and $\mathcal{B}$. In practice, $L=NK$ where $K$ is the number of training examples for each class. By concatenating these two sequences, we obtain one long sequence representing continual learning examples to be presented as an input sequence to a (left-to-right) auto-regressive model. At the end of the sequence, the model is tasked to make predictions on test examples sampled from both $\mathcal{A^{\prime}}$ and $\mathcal{B^{\prime}}$; we assume a single test example for each task (hence, without index): $({\bm{x}}^{\mathcal{A^{\prime}}},y^{\mathcal{A^{\prime}}})$ and $({\bm{x}}^{\mathcal{B^{\prime}}},y^{\mathcal{B^{\prime}}})$ respectively; let us further simplify the notation and denote them as $({\bm{x}}^{\mathcal{A}}_{\text{test}},y^{\mathcal{A}}_{\text{test}})$ and $({\bm{x}}^{\mathcal{B}}_{\text{test}},y^{\mathcal{B}}_{\text{test}})$ instead.
@@
Our model is a self-referential neural network that modifies its own weight matrices as a function of input observations. To simplify the notation, we denote the state of our self-referential NN as a single SRWM ${\bm{W}}_{*}$ (even though, in practice, it may have many of them) where we’ll replace *** by various symbols representing the context/inputs it has observed. Given a training sequence $\left(({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L}),({\bm{x}}^{\mathcal{B}}_{1},y^{\mathcal{B}}_{1}),...,({\bm{x}}^{\mathcal{B}}_{L},y^{\mathcal{B}}_{L})\right)$, our model consumes one input at a time, from left to right, in the auto-regressive fashion. Let ${\bm{W}}_{\mathcal{A}}$ denote the state of the SRWM that has consumed the first part of the sequence corresponding to the examples from Task ${\mathbf{A}}$, i.e., $({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L})$, and let ${\bm{W}}_{\mathcal{A},\mathcal{B}}$ denote the state of our SRWM having observed the entire sequence.
@@
ACL Meta-Training Objectives. The ACL objective function consists in tasking the model to correctly predict the test examples of all tasks learned so far at each task boundaries. That is, in the case of two-task scenario described above (learning Task ${\mathbf{A}}$ then Task ${\mathbf{B}}$), we use the weight matrix ${\bm{W}}_{\mathcal{A}}$ to predict the label $y^{\mathcal{A}}_{\text{test}}$ from input $({\bm{x}}^{\mathcal{A}}_{\text{test}},\varnothing)$, and we use the weight matrix ${\bm{W}}_{\mathcal{A},\mathcal{B}}$ to predict the label $y^{\mathcal{B}}_{\text{test}}$ from input $({\bm{x}}^{\mathcal{B}}_{\text{test}},\varnothing)$ as well as the label $y^{\mathcal{A}}_{\text{test}}$ from input $({\bm{x}}^{\mathcal{A}}_{\text{test}},\varnothing)$. By letting $p(y|{\bm{x}};{\bm{W}}_{*})$ denote the model’s output probability for label $y\in\{1,..,N\}$ given input ${\bm{x}}$ and model weights/state ${\bm{W}}_{*}$, the ACL objective function can be expressed as:
@@
$\mathop{\text{missing}}{minimize}_{\theta}-\left(\log(p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A}}))+\log(p(y^{\mathcal{B}}_{\text{test}}|{\bm{x}}^{\mathcal{B}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}}))+\log(p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}}))\right)$ (4)
@@
for an input meta-training sequence $\left(({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L}),({\bm{x}}^{\mathcal{B}}_{1},y^{\mathcal{B}}_{1}),...,({\bm{x}}^{\mathcal{B}}_{L},y^{\mathcal{B}}_{L})\right)$ (extensible to mini-batches with multiple such sequences), where $\theta$ denotes the model parameters (for the SRWM layer, it is the initial weights ${\bm{W}}_{0}$). Figure 1 illustrates the overall meta-training process of ACL.

### v2

@@ Abstract
General-purpose learning systems should improve themselves in open-ended fashion in ever-changing environments. Conventional learning algorithms for neural networks, however, suffer from catastrophic forgetting (CF)—previously acquired skills are forgotten when a new task is learned. Instead of hand-crafting new algorithms for avoiding CF, we propose Automated Continual Learning (ACL) to train self-referential neural networks to meta-learn their own in-context continual (meta-)learning algorithms. ACL encodes all desiderata—good performance on both old and new tasks—into its meta-learning objectives. Our experiments demonstrate that ACL effectively solves “in-context catastrophic forgetting”; our ACL-learned algorithms outperform hand-crafted ones, e.g., on the Split-MNIST benchmark in the replay-free setting, and enables continual learning of diverse tasks consisting of multiple few-shot and standard image classification datasets.footnotetext: Our code is public: https://github.com/IDSIA/automated-cl
in-context learning,continual learning,meta-learning,Transformers
@@ 1 Introduction
Here, instead of hand-crafting learning algorithms for continual learning, we train self-referential neural networks (Schmidhuber, 1992a, 1987) to meta-learn their own “in-context” continual learning algorithms. We train them through gradient descent on learning objectives that reflect desiderata for continual learning algorithms—good performance on both old and new tasks, including forward and backward transfer. In fact, by extending the standard settings of few-shot or meta-learning based on sequence-processing NNs (Hochreiter et al., 2001; Younger et al., 1999; Cotter & Conwell, 1991, 1990; Mishra et al., 2018), the continual learning problem can also be formulated as a long-span sequence processing task (Irie et al., 2022c). Corresponding CL sequences can be obtained by concatenating multiple few-shot/meta-learning sub-sequences, where each sub-sequence consists of input/target examples corresponding to the task to be learned in-context. As we’ll see in Sec. 3, this setting also allows us to seamlessly express classic desiderata for CL as part of objective functions of the meta-learner.
Once formulated as such a sequence-learning task, we let gradient descent search for CL algorithms achieving the desired CL behaviors in the program space of NN weights. In principle, all typical challenges of CL—such as the stability-plasticity dilemma (Grossberg, 1982)—are automatically discovered and handled by the gradient-based program search process. Once trained, CL is automated through recursive self-modification dynamics of the trained NN, without requiring any human intervention such as adding extra regularization or setting hyper-parameters for CL. Therefore, we call our method, Automated Continual Learning (ACL).
@@ 2 Background
Here we briefly review continual learning and its desiderata (Sec. 2.1), few-shot/meta-learning via sequence processing (Sec. 2.2), and self-modifying NNs (Sec. 2.3), that are essential concepts for describing our method in Sec. 3.
@@ 2.1 Continual Learning
Continual or lifelong learning is a special form of multi-task learning (Thrun, 1998; Caruana, 1997). In conventional multi-task learning scenarios, all datasets for different tasks are available at once, and NNs are trained jointly on all of them (e.g., training batches mix examples from different datasets without any particular order). In contrast, in CL settings, presentation of tasks/datasets is sequential; an NN training process only has access to one task/dataset at a time. When a dataset for a new task becomes available, we lose access to the training examples from the previous task.
@@
Classic desiderata for a CL system (see, e.g., Lopez-Paz & Ranzato (2017); Veniat et al. (2021)) are typically summarized as good performance on three metrics: classification accuracies on each dataset (their average), backward transfer (i.e., impact of learning a new task on the model’s performance on previous tasks; e.g., catastrophic forgetting is a negative backward transfer), and forward transfer (impact of learning a task for the model’s performance on a future task). From a broader perspective of meta-learning systems, we may also measure other effects such as learning acceleration (i.e., whether the system leverages previous learning experiences to accelerate future learning); here our primary focus remains the classic CL metrics above.
@@ 2.2 Few-shot/meta-learning via Sequence Learning
In Sec. 3, we’ll formulate continual learning as a long-span sequence processing task. This is a direct extension of the classic few-shot/meta learning formulated as a sequence learning problem. In fact, since the seminal works by Hochreiter et al. (2001); Younger et al. (1999); Cotter & Conwell (1991, 1990) (see also Naik & Mammone (1992)), many sequence processing neural networks (see, e.g., Bosc (2015); Santoro et al. (2016); Duan et al. (2016); Wang et al. (2017); Munkhdalai & Yu (2017); Munkhdalai & Trischler (2018); Miconi et al. (2018, 2019); Munkhdalai et al. (2019); Kirsch & Schmidhuber (2021); Sandler et al. (2021); Huisman et al. (2023) including “quadratic” Transformers (Vaswani et al., 2017) as in Mishra et al. (2018)) have been trained as a meta-learner (Schmidhuber, 1987, 1992a) that learn by observing sequences of training examples (i.e., pairs of inputs and their labels) in-context. Here we briefly review such a formulation.
@@
Let $d$, $N$, $K$, $P$ be positive integers. In sequential $N$-way $K$-shot classification settings, a sequence processing NN with a parameter vector $\theta\in\mathbb{R}^{P}$ observes a pair (${\bm{x}}_{t}$, $y_{t}$) where ${\bm{x}}_{t}\in\mathbb{R}^{d}$ is the input and $y_{t}\in\{1,...,N\}$ is its label at each step $t\in\{1,...,N\cdot K\}$, corresponding to $K$ examples for each one of $N$ classes. After the presentation of these $N\cdot K$ examples (often called the support set), one extra input ${\bm{x}}\in\mathbb{R}^{d}$ (often called the query) is fed to the model without its true label but with an “unknown label” token $\varnothing$ (number of input labels accepted by the model is thus $N+1$). The model is trained to predict its true label, i.e., the parameters of the model $\theta$ are optimized to maximize the probability $p(y|({\bm{x}}_{1},y_{1}),...,({\bm{x}}_{N\cdot K},y_{N\cdot K}),({\bm{x}},\varnothing);\theta)$ of the correct label $y\in\{1,...,N\}$ of the input query ${\bm{x}}$. Since class-to-label associations are randomized and unique to each sequence ($({\bm{x}}_{1},y_{1}),...,({\bm{x}}_{N\cdot K},y_{N\cdot K}),({\bm{x}},\varnothing)$), each such a sequence represents a new (few-shot or meta) learning example to train the model. To be more specific, this is the synchronous label setting of Mishra et al. (2018) where the learning phase (observing examples, $({\bm{x}}_{1},y_{1})$ etc.) is separated from the prediction phase (predicting label $y$ given $({\bm{x}},\varnothing)$). We opt for this variant in our experiments as we empirically find this (at least in our specific settings) more stable than the delayed label setting (Hochreiter et al., 2001) where the model has to make a prediction for every input, and the label is fed to the model with a delay of one time step.
2.3 Self-Referential Weight Matrices
@@
Our method (Sec. 3) can be applied to any sequence-processing NN architectures in principle. Nevertheless, certain architectures naturally fit better to parameterize a self-improving continual learner. Here we use the modern self-referential weight matrix (SRWM; Irie et al. (2022c, 2023)) to build a generic self-modifying NN. An SRWM is a weight matrix that sequentially modifies itself as a response to a stream of input observations (Schmidhuber, 1992a, 1993). The modern SRWM belongs to the family of linear Transformers (LTs) a.k.a. Fast Weight Programmers (FWPs; Schmidhuber (1991, 1992b); Katharopoulos et al. (2020); Choromanski et al. (2021); Peng et al. (2021); Schlag et al. (2021); Irie et al. (2021a)). Linear Transformers and FWPs are an important class of the now popular Transformers (Vaswani et al., 2017): unlike the standard ones whose computational requirements grow quadratically and whose state size grows linearly with the context length, LTs/FWPs’ complexity is linear and the state size is constant w.r.t. sequence length (like in the standard RNNs). This is an important property for in-context CL, since, conceptually, we want such a CL system to continue to learn for an arbitrarily long, lifelong time span. Moreover, the duality between linear attention and FWPs (Schlag et al., 2021)—and likewise, between linear attention and gradient descent-trained linear layers (Irie et al., 2022a; Aizerman et al., 1964)—have played a key role in certain theoretical analyses of in-context learning capabilities of Transformers (von Oswald et al., 2023a; Dai et al., 2023).
@@
The dynamics of an SRWM (Irie et al., 2022c) are described as follows. Let $d_{\text{in}}$, $d_{\text{out}}$, $t$ be positive integers, and $\otimes$ denote outer product. At each time step $t$, an SRWM ${\bm{W}}_{t-1}\in\mathbb{R}^{(d_{\text{out}}+2*d_{\text{in}}+1)\times d_{\text{in}}}$ observes an input ${\bm{x}}_{t}\in\mathbb{R}^{d_{\text{in}}}$, and outputs ${\bm{y}}_{t}\in\mathbb{R}^{d_{\text{out}}}$, while also updating itself to ${\bm{W}}_{t}$ as:
@@
where ${\bm{v}}_{t},\bar{{\bm{v}}}_{t}\in\mathbb{R}^{(d_{\text{out}}+2*d_{\text{in}}+1)}$ are value vectors, ${\bm{q}}_{t}\in\mathbb{R}^{d_{\text{in}}}$ and ${\bm{k}}_{t}\in\mathbb{R}^{d_{\text{in}}}$ are query and key vectors, and $\sigma(\beta_{t})\in\mathbb{R}$ is the learning rate. $\sigma$ and $\phi$ denote sigmoid and softmax functions respectively. $\phi$ is typically also applied to ${\bm{x}}_{t}$ in Eq. 1; here we follow Irie et al. (2022c)’s few-shot image classification setting, and use the variant without it. Eq. 3 corresponds to a rank-one update of the SRWM, from ${\bm{W}}_{t-1}$ to ${\bm{W}}_{t}$, through the delta learning rule (Widrow & Hoff, 1960; Schlag et al., 2021) where the self-generated patterns, ${\bm{v}}_{t}$, $\phi({\bm{k}}_{t})$, and $\sigma(\beta_{t})$, play the role of target, input, and learning rate of the learning rule respectively. The delta rule is crucial for the performance of LTs (Schlag et al., 2021; Irie et al., 2021a, 2022b; Irie & Schmidhuber, 2023b).
The initial weight matrix ${\bm{W}}_{0}$ is the only trainable parameters of this layer, that encodes the initial self-modification algorithm. In practice, we use the layer above as a direct replacement to the self-attention layer in the Transformer architecture (Vaswani et al., 2017); we also use the multi-head version of the SRWM computation above (Irie et al., 2022c).
@@ 3 Method
Task Formulation. We formulate continual learning as a long-span sequence learning task. Let $D$, $N$, $K$, $L$ denote positive integers. Consider two $N$-way classification tasks ${\mathbf{A}}$ and ${\mathbf{B}}$ to be learned sequentially (as we’ll see, this can be straightforwardly extended to more tasks). The formulation here applies to both “meta-training” and “meta-test” phases (see Appendix A.1 for more on this terminology). We denote the respective training datasets as $\mathcal{A}$ and $\mathcal{B}$, and test sets as $\mathcal{A^{\prime}}$ and $\mathcal{B^{\prime}}$. We assume that each datapoint in these datasets consists of one input feature ${\bm{x}}\in\mathbb{R}^{D}$ of dimension $D$ (generically denoted as vector ${\bm{x}}$, but it is an image in all our experiments) and one label $y\in\{1,...,N\}$. We consider two sequences of $L$ training examples $\left(({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L})\right)$ and $\left(({\bm{x}}^{\mathcal{B}}_{1},y^{\mathcal{B}}_{1}),...,({\bm{x}}^{\mathcal{B}}_{L},y^{\mathcal{B}}_{L})\right)$ sampled from the respective training sets $\mathcal{A}$ and $\mathcal{B}$. In practice, $L=NK$ where $K$ is the number of training examples for each class. By concatenating these two sequences, we obtain one long sequence representing CL examples to be presented as an input sequence to a (left-to-right) auto-regressive model. At the end of the sequence, the model is tasked to make predictions on test examples sampled from both $\mathcal{A^{\prime}}$ and $\mathcal{B^{\prime}}$; we assume a single test example for each task (hence, without index): $({\bm{x}}^{\mathcal{A^{\prime}}},y^{\mathcal{A^{\prime}}})$ and $({\bm{x}}^{\mathcal{B^{\prime}}},y^{\mathcal{B^{\prime}}})$ respectively; which we simply denote as $({\bm{x}}^{\mathcal{A}}_{\text{test}},y^{\mathcal{A}}_{\text{test}})$ and $({\bm{x}}^{\mathcal{B}}_{\text{test}},y^{\mathcal{B}}_{\text{test}})$ instead.
@@
Our model is a self-referential NN that modifies its own weight matrices as a function of input observations. To simplify the notation, we denote the state of our self-referential NN as a single SRWM ${\bm{W}}_{*}$ (even though, in practice, it may have many of them) where we’ll replace * by various symbols representing the context/inputs it has observed. Given a training sequence $\left(({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L}),({\bm{x}}^{\mathcal{B}}_{1},y^{\mathcal{B}}_{1}),...,({\bm{x}}^{\mathcal{B}}_{L},y^{\mathcal{B}}_{L})\right)$, our model auto-regressively consumes one input at a time, from left to right, in the auto-regressive fashion. Let ${\bm{W}}_{\mathcal{A}}$ denote the state of the SRWM that has consumed the first part of the sequence corresponding to the examples from Task ${\mathbf{A}}$, i.e., $({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L})$, and let ${\bm{W}}_{\mathcal{A},\mathcal{B}}$ denote the state of our SRWM having observed the entire sequence.
@@
ACL Meta-Training Objectives. The ACL meta-training objective function tasks the model to correctly predict the test examples of all tasks learned so far at each task boundaries. That is, in the case of two-task scenario described above (learning Task ${\mathbf{A}}$ then Task ${\mathbf{B}}$), we use the weight matrix ${\bm{W}}_{\mathcal{A}}$ to predict the label $y^{\mathcal{A}}_{\text{test}}$ from input $({\bm{x}}^{\mathcal{A}}_{\text{test}},\varnothing)$, and we use the weight matrix ${\bm{W}}_{\mathcal{A},\mathcal{B}}$ to predict the label $y^{\mathcal{B}}_{\text{test}}$ from input $({\bm{x}}^{\mathcal{B}}_{\text{test}},\varnothing)$ as well as the label $y^{\mathcal{A}}_{\text{test}}$ from input $({\bm{x}}^{\mathcal{A}}_{\text{test}},\varnothing)$. By letting $p(y|{\bm{x}};{\bm{W}}_{*})$ denote the model’s output probability for label $y\in\{1,..,N\}$ given input ${\bm{x}}$ and model weights/state ${\bm{W}}_{*}$, the ACL objective can be expressed as:
@@
$\mathop{\text{missing}}{minimize}_{\theta}\;-\Bigl{(}\log(p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A}}))\;+$ (4)
$\;\;\log(p(y^{\mathcal{B}}_{\text{test}}|{\bm{x}}^{\mathcal{B}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}}))+\log(p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}}))\Bigr{)}$
@@
for an arbitrary input meta-training sequence $\left(({\bm{x}}^{\mathcal{A}}_{1},y^{\mathcal{A}}_{1}),...,({\bm{x}}^{\mathcal{A}}_{L},y^{\mathcal{A}}_{L}),({\bm{x}}^{\mathcal{B}}_{1},y^{\mathcal{B}}_{1}),...,({\bm{x}}^{\mathcal{B}}_{L},y^{\mathcal{B}}_{L})\right)$ (which is extensible to mini-batches with multiple such sequences), where $\theta$ denotes the model parameters (for the SRWM layer, it is the initial weights ${\bm{W}}_{0}$). Figure 1 illustrates the overall meta-training process of ACL.